how do solve this 4y(7y - 6) + 20y
Answer by richwmiller(17219)

Before we solve your problem.
What would you do with 2*(3+4)
You can add the 3and 4 first to get 7 and then multiply by 2 and get 14
or you can go 2*6+2*4=8+6=14
In your problem we are doing it thie second way because we can't do it the first way.
4y(7y - 6) + 20y
4y*7y-4y*(-6)+20y
remember -4y*-6 will be positive minus times minus is plus
continue